OpenStudy (jolo_yolo):

Choose the equivalent system of linear equations that will produce the same solution as the one given below. 6x + 2y = -6 3x - 4y = -18 12x + 4y = -12 15x = -30 8x + 4y = -4 14x = -10 6x - 8y = -36 -6y = -42 6x - y = 15 3y = -9

OpenStudy (anonymous):

i thought there might be some snappy way to do this, but down i do not see it guess we need to solve \[6x+2y=-6\\ 3x-4y=-18\] and get \[x=-2,y=3\] and check against the others

OpenStudy (anonymous):

fortunately we do not have too far to check , since it works for the first one \[12x+4y=-12\\ 15x=-30\]

OpenStudy (jolo_yolo):

oh okay so the answer turns around to A